OpenStudy (anonymous):

Use Direct Comparison to determine whether the infinite series 1/n! converges or diverges

OpenStudy (anaise):

Bruh...

OpenStudy (anonymous):

Notice that \[\frac{1}{n!}=\frac{1}{n(n-1)\cdots2\times1}\le\frac{1}{n(n-1)}\le\frac{1}{n^2}\] What do you know about the series \(\displaystyle\sum_{n\ge1}\frac{1}{n^2}\)? Alternatively, if you've heard of power series, then you've no doubt become at least somewhat familiar with the exponential function's series representation.